OpenStudy (nykoreyan):

Find a solution to the following system of equations. -5x+y=-5 -4x+2y=2 a.(–2, –15) b.(2, 5) c.(0, 1) d.(–8, –15)

OpenStudy (anonymous):

B. (2,5).

OpenStudy (anonymous):

There's a cool graphing calculator site. Just put in the equations, it graphs it. I like it a lot. You should check it out. It's easy to use. www.desmos.com Go there, then there's a red button in the top right corner of the webpage.

OpenStudy (anonymous):

still need help?

OpenStudy (anonymous):

Problem: Solve −5x+y=−5;−4x+2y=2 Steps: I will try to solve your system of equations. −5x+y=−5;−4x+2y=2 Step: Solve−5x+y=−5for y: −5x+y+5x=−5+5x(Add 5x to both sides) y=5x−5 Step: Substitute5x−5foryin−4x+2y=2: −4x+2y=2 −4x+2(5x−5)=2 6x−10=2(Simplify both sides of the equation) 6x−10+10=2+10(Add 10 to both sides) 6x=12 6x6=126(Divide both sides by 6) x=2 Step: Substitute2forxiny=5x−5: y=5x−5 y=(5)(2)−5 y=5(Simplify both sides of the equation) Answer: y=5 and x=2

OpenStudy (nykoreyan):

@JasonSpade ik desmos i just dont no how to get the right answer bc its a graph and id have to constantly zoom in

OpenStudy (anonymous):

Just go to where the lines intersect and click it, it will give you the point.